A pie-shaped (triangle) lake-front lot has a perimeter of 900 feet. One side is 100 feet longer than the shortest side,while the third side is 500 feet longer than the shortest side. Find the lengths of all three sides.
Answer by checkley77(12569)

x+(x+100)+x+500)=900
3x=900-600
3x=300
x=100 ft. is the shortest side.
100+100=200 ft. for the next side.
100+500=600 ft. for the longest side.
Proof:
100+200+600=900
900=900